Is there an easier way to solve for x in this equation:
20% (x/0.2 + 3 2/3 - 4.32) = 1/5 (1 2/3 + 2 - 2 - 2 32/100)

Thanks!

Found 2 solutions by Alan3354, biboyjimenez:
Answer by Alan3354(69443)

Easier than what?

Answer by biboyjimenez(1)

20% is similar to 1/5 so they both can be cancelled
x/0.2 + 3 2/3 - 4.32 = 3 2/3 - 2 - 2 32/100
x/0.2 -4.32 = -4.32
x/0.2 = 0
x = 0